OpenStudy (anonymous):

cos^4(3x) - sin^4(3x)

OpenStudy (anonymous):

You need to use the following: \[ \begin{array}{rcl} a^2-b^2 &=& (a-b)(a+b) &\text{Difference of squares} \\ 1 &=& \sin^2(x)+\cos^2(x) &\text{Pythagorean Theorem} \\ \cos(2x) &=& \cos^2(x) - \sin^2(x)&\text{Double angle formula} \end{array} \]